# Uncovering flow and deformation regimes in the coupled fluid-solid vestibular system

### v1

@@ 5.3 Non-uniform channel widths
As noted in the introduction, an advantage of our theoretical formulation is that it is also compatible with a non-uniform and arbitrary channel width, described by the function $a(s)$, so long as the small aspect ratio between channel width and length is maintained. This case is more delicate than the one we saw in the last subsection, as the pressure gradient is no longer constant but depends on $a(s)$, and must be integrated along the channel to obtain the pressure jump across the cupula. In Appendix C a method based on the Laplace transform and the convolution theorem is developed, through which we obtain an approach to solving the problem for both non-uniform channel widths and $\mathrm{St}>0$. The main result is that we obtain the same form of equation (42) for the deflection $\eta_{0}(r,t)$, with the Kernel $\mathcal{K}(x;\mathrm{St})$ given by the (temporal) inverse Laplace transform of
$\tilde{\mathcal{K}}(\sigma;\mathrm{St})=\left(\int_{0}^{2\pi}\frac{\mathrm{d}s}{a(s)^{4}\sum_{n=0}^{\infty}\lambda_{n}^{-2}\left[a(s)^{2}\mathrm{St}\sigma+\lambda_{n}^{2}\right]^{-1}}\right)^{-1}.$ (51)
@@ 5.4 Fluid inertia explains discrepancies between numerics and model
Figure 4 generally shows excellent agreement between the $\mathrm{St}=0$ model presented in § 3 and our COMSOL numerics, especially at the scale of the largest velocities, which were used for comparison in figure 4. However, if we zoom in to situations where the velocity is small, such as when $t=0.5$, then we might expect to observe differences caused by small errors in the phase of the motion. Figure 9 shows just such an effect: the agreement between the predictions of the $\mathrm{St}=0$ asymptotics (dashed black curves) and COMSOL simulations (blue markers) is no longer satisfactory for small and moderate values of the stiffness $\kappa\lesssim 1$: while the absolute error is small, the relative error is very large. This is because even for small Stokes numbers, the exact time at which the velocity zero is different from that predicted by the $\mathrm{St}=0$ asymptotics of § 3, which essentially assume that the motion is quasi-steady.
Figure 9 also shows the predictions of the leading-order in $\mathrm{St}$ asymptotic results presented in this section. As might be expected, we see much better agreement between the prediction accounting for $\mathrm{St}>0$ via (38b) and (43) (shown by solid red curves in figure 9) and the results of COMSOL simulations (points) than with the earlier result, which neglected the effects of fluid inertia entirely. We emphasize that this finite inertia case requires the numerical calculation of the integrals in (38b) and (43).
Figure 9: Velocity profiles for $t=0.5$, including the finite fluid inertia correction. Dashed black line: $\mathrm{St}=0$, blue markers: numerics, solid red line: $\mathrm{St}>0$ model.
@@ 6 Analytical description of vortical flow
Mathematically, the reason why we might expect enlarged regions of the canal to experience symmetry breaking may be seen from the form of the leading order and $O(\epsilon)$ axial velocity terms in (29). In particular, in regions where $a(s)>1$, the magnitude of the leading order velocity $w_{0}$ is proportional to $1/[I_{4}\,a(s)^{4}]$, while the correction is of order $\epsilon$ the $f(t)$ term in (29b) remains $O(\epsilon)$ even as $a(s)^{4}$ becomes large. (We retain the $I_{4}$ term here, since it is also dependent on $a(s)$.) Hence, we might expect noticeable asymmetry in the flow to develop in regions where $1/[I_{4}a(s)^{4}]$ is comparable to $\epsilon$. Note that this can be achieved in this way without requiring that we are in the stiff cupula regime $\kappa\gg 1$ that led to the symmetry-breaking discussed in Section 3.2.
@@
$a(s)=1+(a_{m}-1)e^{-\gamma(s-\pi)^{2}}.$ (52)
Here, the cupula and utricle are located at $s=\pi$, $\gamma$ is a parameter controlling the width of the enlargement and $a_{m}$ is the maximum inner radius of the tube. This choice is motivated by the qualitative agreement with the imaging from Daocai et al. (2014). In Figure 10 we plot the velocity distribution for several channel geometries, in particular a top view of the mid plane of the flow around the canal, with colour indicating the magnitude of the axial velocity. The size of the enlarged region is increased from left to right (with $\gamma=1$ and $a_{m}=1,2,3,4$), and the appearance of the vortical flow is clear. The forcing is given by (34) and we visualize the solution at time $t=0.25$. Here we have used a small value of $\kappa=0.1$ so that the flow in the slender regions is predicted to remain largely symmetric, a feature that we will verify below.
@@ 6.1 Conditions required for the formation of the utricular vortex
$|w_{0}^{*}(s,t)|=\frac{\pi}{2I_{4}a(s)^{2}}\left|(t)+\frac{\kappa\Delta p_{0}}{2\pi}\right|\sim\frac{\pi}{2I_{4}a(s)^{2}}|(f(t)|),$ (53)
@@
$|w_{1}^{*}|=\frac{1}{24\sqrt{3}I_{4}a(s)}\left|-3\kappa\Delta p_{0}+\left(-6\pi+4I_{4}a(s)^{4}\right)f(t)\right|.$ (54)
Focusing on the utricle, where $a(s)$ is largest, $|w_{1}^{*}|$ is dominated by
$|w_{1}^{*}|\sim\frac{a_{m}^{3}}{6\sqrt{3}}\left|f(t)\right|.$ (55)
@@
The transition can be seen qualitatively in the left panel of Figure 10, where the flow in the utricle (left half of the torus) transitions from symmetric to asymmetric as the bulge is increased. In Figure 11 this transition is analysed quantitatively. In the left panel, the flow profile in the utricle is plotted for varying $a_{m}$. Observe that for large $a_{m}$, the flow utricle flow is asymmetric while the flow in the slender part of the canal remains symmetric. To quantify the transition to noticeably asymmetrical flow, we compute the correlation between the velocity profile $w=w_{0}+\epsilon w_{1}$ and the symmetric ($w_{0}$) and asymmetric ($w_{1}$) solutions. The correlations may be written as
$R(w,w_{0})=\int_{0}^{a}w(r)w_{0}(r)\mathrm{d}r=\sqrt{\frac{R_{0}}{R_{0}+\epsilon^{2}R_{1}}},$ (57a)
$R(w,w_{1})=\int_{0}^{a}w(r)w_{1}(r)\mathrm{d}r=\sqrt{\frac{R_{1}}{R_{0}+\epsilon^{2}R_{1}}},$ (57b)
$R_{0}=\int_{-a}^{a}w_{0}^{2}\mathrm{d}r=\frac{1}{15a_{m}^{3}I_{4}^{2}}\left(\kappa\Delta p_{0}+2\pi f(t)\right)^{2},$ (57c)
@@
$R_{1}=\int_{-a}^{a}w_{1}^{2}\mathrm{d}r=\frac{1}{1680I_{4}^{2}a_{m}}\left(3\pi\Delta p_{0}+(6\pi-4I_{4}a_{m}^{4})f(t)\right).$ (57d)
@@
Here, we have used the fact that $\int_{-a}^{a}w_{0}w_{1}\mathrm{d}r=0$, which is trivial since $w_{0}w_{1}$ is an odd function of $r$. The correlation $R(w,w_{0})$ will be close to 1 when the flow is largely symmetrical, and closer to zero when the flow is noticeably asymmetrical. The opposite is true of the correlation $R(w,w_{1})$. In the the right panel of fig. 11 we plot the symmetrical correlation $R(w,w_{0})$ (solid line) and the asymmetrical correlation $R(w,w_{1})$ (dashed line) as predicted from our analytical model.
@@
Based on the definition of $\xi$ we expect a transition when $\xi\sim 1$. Fig. 11 confirms this expectation, showing that a transition indeed occurs at this point when the geometry $a(s)$ satisfies (52). We note that for times when $f(t)=0$, our analysis does not apply, and the flow remains symmetric, even for large utricles. This follows from (55), as it is clear that $w_{1}=0$ when $f(t)=0$.
@@
Our analysis of the onset of vortical flow is further supported by numerical simulations, represented by triangles and stars in the righthand panel of fig. 11. These simulations were conducted using COMSOL, modelling the utricle as an ellipsoidal expansion of the toroidal geometry. We find good agreement between the analytical predictions and COMSOL simulations regarding the emergence of vortical flow. (We attribute the small discrepancy between the COMSOL and analytical model to the fact that the geometry is not identical in both cases.) Furthermore, the consistency between the numerical simulations and analytical solutions suggests that, although the first-order correction may be larger than the leading-order term, the subsequent terms in the series remain well behaved, ensuring that the asymptotic ordering is preserved.
@@
Figure 11: Left panel: analytical reconstruction of flow profiles in the wide region of the channel (representing the utricle), $w(r,\theta,s=\pi,t)$ for different values of the maximum enlargement $a_{m}$. The inset shows the flow profiles in the thin region of the flow $w(r,\theta,s=0,t)$; these remain symmetric, confirming the symmetry breaking mechanism is not the same as the global symmetry-breaking mechanism discussed in § 3.2. Right panel: correlation (as defined in (57)) between the axial velocity in the utricle $w(r)=w_{0}(r)+\epsilon w_{1}(r,\theta)$ and the symmetric (solid) and asymmetric flow profile (dashed). Curves show the results of the analytical computation and triangles and stars show the correlations computed from the COMSOL solution. We find that the transition occurs when $\xi=a_{m}^{5}I_{4}\epsilon/(3\sqrt{3}\pi)\sim 1$, as predicted by our analysis.
@@ 7 Discussion and conclusion
Our analytical approach consisted in solving the Navier-Stokes equations via an asymptotic series in the small aspect ratio of the semicircular canals. Through asymptotic analysis, and by connecting the fluid flow at the cupula to the cupular deformation, described via a plate equation, we reduced the vestibular dynamics to an ODE system for the cupular deflection, whose behaviour could easily be characterized. In this way, we established two primary regimes of flow-cupula interaction, depending on the value of the relative stiffness parameter $\kappa$:
@@
  • Soft cupula regime, $\kappa\ll 1$: When the cupular stiffness is relatively low, the deformation of the cupula closely follows the angular velocity of the head. In this regime, the flow in the canal exhibits symmetry about the centreline.
@@
  • Stiff cupula regime, $\kappa\gg 1$: As the stiffness of the cupula increases, the deformation transitions to follow the angular acceleration of the head. In this regime, the symmetry of the fluid flow about the centreline is broken, creating distinct zones of differential flow. This transition highlights the importance of structural properties of the cupula in shaping the dynamic response of the vestibular system.
@@
To verify the analytical findings, we conducted numerical simulations of the reduced equations using COMSOL Multiphysics. The numerical results showed excellent agreement with the asymptotic predictions, confirming the validity of the analytical approximations across a wide parameter space. Importantly, the numerical approach enabled us to explore the influence of complex fluid–solid interaction boundary conditions and the nonlinear, advection, term in the Navier-Stokes equations that are otherwiseintractable analytically.
@@
When the inertial terms were incorporated into the governing equations via inclusion of a finite Stokes number, $\mathrm{St}$, we observed important modifications to the system’s behaviour. For small Stokes numbers, and in studying the relaxation of the cupula to an initial deformation, the system exhibited overdamped dynamics. This behaviour is also predicted when inertia is neglected, and is consistent with the low-Reynolds-number assumption inherent to the vestibular fluid dynamics. However, for sufficiently large Stokes numbers, the system exhibited underdamped oscillations, with the cupular deformation following angular acceleration even in the soft cupula regime. This transition highlights the interplay between inertial and viscous forces in shaping the dynamic response of the system. Physiologically, this finding suggests that under certain conditions, such as during rapid head movements, the vestibular system may exhibit enhanced sensitivity to acceleration due to inertial effects.
@@
In each of the scenarios considered, our analytical approach enabled us to derive explicit expressions for the transition point between physical regimes, that is we obtained formulas for the critical values of the relevant system parameters at which the transition between different regimes occurs. These formulas lend insight into the fine balance between different components of the system, and enable to speculate on how the vestibular system may have been fine tuned by evolution in different organisms, and/or key considerations in engineering an artificial vestibular system. We turn to such considerations next.

### v2

@@ 5.3 Non-uniform channel widths
As noted in the introduction, an advantage of our theoretical formulation is that it is also compatible with a non-uniform and arbitrary channel width, described by the function $a(s)$, so long as the small aspect ratio between channel width and length is maintained. This case is more delicate than the one we saw in the last subsection, as the pressure gradient is no longer constant but depends on $a(s)$, and must be integrated along the channel to obtain the pressure jump across the cupula. In Appendix D a method based on the Laplace transform and the convolution theorem is developed, through which we obtain an approach to solving the problem for both non-uniform channel widths and $\mathrm{St}>0$. The main result is that we obtain the same form of equation (48) for the deflection $\eta_{0}(r,t)$, with the kernel $\mathcal{K}(x;\mathrm{St})$ given by the (temporal) inverse Laplace transform of
$\tilde{\mathcal{K}}(\sigma;\mathrm{St})=2\pi\left(\int_{0}^{2\pi}\frac{\mathrm{d}s}{a(s)^{4}\sum_{n=1}^{\infty}\lambda_{n}^{-2}\left[a(s)^{2}\mathrm{St}\sigma+\lambda_{n}^{2}\right]^{-1}}\right)^{-1}.$ (57)
@@ 5.4 Fluid inertia explains discrepancies between numerics and model
Figure 4 generally shows excellent agreement between the $\mathrm{St}=0$ model presented in §3 and our COMSOL numerics, especially at the scale of the largest velocities, which were used for comparison in figure 4. However, if we zoom in to situations where the velocity is small, such as when $t=0.5$, then we might expect to observe differences caused by small errors in the phase of the motion. Figure 9 shows just such an effect: the agreement between the predictions of the $\mathrm{St}=0$ asymptotics (dashed black curves) and COMSOL simulations (blue markers) is no longer satisfactory for small and moderate values of the stiffness $\kappa\lesssim 1$: while the absolute error is small, the relative error is very large. This is because even for small Stokes numbers, the exact time at which the velocity vanishes is different from that predicted by the $\mathrm{St}=0$ asymptotics of §3, which essentially assume that the motion is quasi-steady.
Figure 9 also shows the predictions of the leading-order in $\mathrm{St}$ asymptotic results presented in this section. As might be expected, we see much better agreement between the prediction accounting for $\mathrm{St}>0$ via (43d) and (48) (shown by solid red curves in figure 9) and the results of COMSOL simulations (points) than with the earlier result, which neglected the effects of fluid inertia entirely. We emphasize that this finite inertia case requires the numerical calculation of the integrals in (43d) and (48).
Figure 9: Velocity profiles for $t=0.5$, including the finite fluid inertia correction. Dashed black line: $\mathrm{St}=0$, blue markers: numerics, solid red line: $\mathrm{St}>0$ model. The parameter values are the same as in figure 7.
@@ 6 Analytical description of vortical flow
Mathematically, the reason why we might expect enlarged regions of the canal to experience symmetry breaking may be seen from the form of the leading order and $O(\epsilon)$ axial velocity terms in (34). In particular, in regions where $a(s)>1$, the magnitude of the leading order velocity $w_{0}$ is proportional to $1/[I_{4}\,a(s)^{4}]$, while the correction is of order $\epsilon$ the $\dot{\Omega}(t)$ term in (34b) remains $O(\epsilon)$ even as $a(s)^{4}$ becomes large. (We retain the $I_{4}$ term here, since it is also dependent on $a(s)$.) Hence, we might expect noticeable asymmetry in the flow to develop in regions where $1/[I_{4}a(s)^{4}]$ is comparable to $\epsilon$. Note that this can be achieved in this way without requiring that we are in the stiff cupula regime $\kappa\gg 1$ that led to the symmetry-breaking discussed in Section 3.2.
@@
$a(s)=1+(a_{m}-1)e^{-\gamma s^{2}}.$ (58)
Here, the cupula and utricle are located at $s=0$, $\gamma$ is a parameter controlling the width of the enlargement and $a_{m}$ is the maximum inner radius of the tube. This choice is motivated by the qualitative agreement with the imaging from Daocai et al. (2014). In Figure 10 we plot the velocity distribution for several channel geometries, in particular a top view of the mid plane of the flow around the canal, with colour indicating the magnitude of the axial velocity. The size of the enlarged region is increased from left to right (with $\gamma=1$ and $a_{m}=1,2,3,4$), and the appearance of the vortical flow is clear. The forcing is given by (39) and we visualize the solution at time $t=0.25$. Here we have used a small value of $\kappa=0.1$ so that the flow in the slender regions is predicted to remain largely symmetric, a feature that we will verify below.
@@ 6.1 Conditions required for the formation of the utricular vortex
$|w_{0}^{*}(s,t)|=\frac{\pi}{2I_{4}a(s)^{2}}\left|\dot{\Omega}(t)+\frac{\Delta p_{0}}{2\pi}\right|\sim\frac{\pi}{2I_{4}a(s)^{2}}|\dot{\Omega}(t)|,$ (59)
@@
$|w_{1}^{*}|=\frac{1}{24\sqrt{3}I_{4}a(s)}\left|-3\Delta p_{0}+\left(-6\pi+4I_{4}a(s)^{4}\right)\dot{\Omega}(t)\right|.$ (60)
The symmetric component of $w_{1}$ scales as $\epsilon\Delta p_{1}^{\text{outer}}/a(s)^{2}$, and since it is much smaller than both the asymmetric component of $w_{1}$ and the leading-order flow $w_{0}$, it is neglected in the present analysis (we set $\Delta p_{1}^{\text{outer}}=0$). Focusing on the utricle, where $a(s)$ is largest, $|w_{1}^{*}|$ is dominated by
$|w_{1}^{*}|\sim\frac{a_{m}^{3}}{6\sqrt{3}}\left|\dot{\Omega}(t)\right|.$ (61)
@@
This motivates the introduction of a transition parameter $\xi=a_{m}^{5}I_{4}\epsilon/(3\sqrt{3}\pi)$. The transition can be seen qualitatively in the left panel of Figure 10, where the flow in the utricle (bottom half of the torus) transitions from symmetric to asymmetric as the bulge is increased. In Figure 11 this transition is analysed quantitatively. In the left panel, the flow profile in the utricle is plotted for varying $a_{m}$. Observe that for large $a_{m}$, the utricle flow is asymmetric while the flow in the slender part of the canal remains symmetric. To quantify the transition to noticeably asymmetrical flow, we compute the correlation between the velocity profile $w=w_{0}+\epsilon w_{1}$ and the symmetric ($w_{0}$) and asymmetric ($w_{1}$) solutions.
The correlations may be written as
$C(w,w_{0})=\frac{R(w,w_{0})}{\sqrt{R(w,w)R(w_{0},w_{0})}},\quad C(w,w_{1})=\frac{R(w,w_{1})}{\sqrt{R(w,w)R(w_{1},w_{1})}},$ (63a)
$R(u(r,\theta),v(r,\theta))=\int_{0}^{2\pi}\int_{0}^{a_{m}}u(r,\theta)v(r,\theta)r~\mathrm{d}r\mathrm{d}\theta,$ (63b)
and we find
$C(w,w_{0})=\sqrt{\frac{R_{0}}{R_{0}+\epsilon^{2}R_{1}}},\quad C(w,w_{1})=\sqrt{\frac{\epsilon^{2}R_{1}}{R_{0}+\epsilon^{2}R_{1}}},$ (64a)
$R_{0}=R(w_{0},w_{0})=2\pi\int_{0}^{a_{m}}w_{0}(r)^{2}r~\mathrm{d}r=\frac{\pi^{3}}{12I_{4}^{2}a_{m}^{2}}\left(\dot{\Omega}(t)+\frac{\Delta p_{0}}{2\pi}\right)^{2},$ (64b)
@@
$\begin{split}&R_{1}=R(w_{1},w_{1})=\int_{0}^{2\pi}\cos^{2}\theta\mathrm{d}\theta\int_{0}^{a_{m}}\left[\frac{r(r^{2}-a_{m}^{2})(-3\Delta p_{0}+(4a_{m}^{4}I_{4}-6\pi)}{16I_{4}a_{m}^{4}}\right]^{2}r~\mathrm{d}r\\ &\quad\,=\frac{\pi}{24\cdot 16^{2}I_{4}^{2}}\left(-3\Delta p_{0}+(4a_{m}^{4}I_{4}-6\pi)\dot{\Omega}(t)\right)^{2}\end{split}$ (64c)
@@
Here, we have assumed that $\Delta p_{1}^{\text{outer}}=0$ and used the fact that $R(w_{0},w_{1})=0$, which is trivial since $w_{0}w_{1}\sim\cos\theta$. The correlation $C(w,w_{0})$ will be close to 1 when the flow is largely symmetrical, and closer to zero when the flow is noticeably asymmetrical. The opposite is true of the correlation $C(w,w_{1})$. In the the right panel of figure 11 we plot the symmetrical correlation $C(w,w_{0})$ (solid line) and the asymmetrical correlation $C(w,w_{1})$ (dashed line) as predicted from our analytical model.
@@
Based on the definition of $\xi$ we expect a transition when $\xi\sim 1$. figure 11 confirms this expectation, showing that a transition indeed occurs at this point when the geometry $a(s)$ satisfies (58). We note that for times when $\dot{\Omega}(t)=0$, our analysis does not apply, and the flow remains symmetric, even for large utricles. This follows from (61), as it is clear that $w_{1}=0$ when $\dot{\Omega}(t)=0$.
@@
Our analysis of the onset of vortical flow is further supported by numerical simulations, represented by triangles and stars in the righthand panel of figure 11. These simulations were conducted using COMSOL, modelling the utricle as an ellipsoidal expansion of the toroidal geometry. We find good agreement between the analytical predictions and COMSOL simulations regarding the emergence of vortical flow. (We attribute the small discrepancy between the COMSOL and analytical model to the fact that the geometry is not identical in both cases, and furthermore because the analytical result does not have a solid obstacle disrupting the flow.) Furthermore, the consistency between the numerical simulations and analytical solutions suggests that, although the first-order correction may be larger than the leading-order term, the subsequent terms in the series remain well behaved, ensuring that the asymptotic ordering is preserved.
[…]
@@
Figure 11: Left panel: analytical reconstruction of flow profiles in the wide region of the channel (representing the utricle), $w(r,\theta,s=\pi,t)$ for different values of the maximum enlargement $a_{m}$. The inset shows the flow profiles in the thin region of the flow $w(r,\theta,s=0,t)$; these remain symmetric, confirming the symmetry breaking mechanism is not the same as the global symmetry-breaking mechanism discussed in §3.2. Right panel: correlation (as defined in (63a)) between the axial velocity in the utricle $w(r)=w_{0}(r)+\epsilon w_{1}(r,\theta)$ and the symmetric (solid) and asymmetric flow profile (dashed). Curves show the results of the analytical computation and triangles and stars show the correlations computed from the COMSOL solution. We find that the transition occurs when $\xi=a_{m}^{5}I_{4}\epsilon/(3\sqrt{3}\pi)\sim 1$, as predicted by our analysis.
@@ 7 Discussion and conclusion
Our analytical approach consisted in solving the Navier-Stokes equations via an asymptotic series in the small aspect ratio of the semicircular canals. Through asymptotic analysis, and by connecting the fluid flow at the cupula to the cupular deformation, we reduced the vestibular dynamics to an ODE system for the cupular pressure jump, whose behaviour could easily be characterized. In this way, we established three primary regimes of flow-cupula interaction, depending on the value of the relative stiffness parameter $\kappa$:
@@
  • Soft cupula regime, $\kappa\ll 1$: When the cupular stiffness is relatively low, the deformation of the cupula closely follows the angular velocity of the head. In this regime, the flow in the canal exhibits symmetry about the centreline, and in dimensional terms the magntiude of the cupular deformation scales as $\hat{\eta}\sim a^{2}R\Omega_{0}/\nu$.
@@
  • Stiff cupula regime, $\kappa\gg 1$: As the stiffness of the cupula increases, the deformation transitions to follow the angular acceleration of the head. In this regime, the symmetry of the fluid flow about the centreline is broken, creating distinct zones of differential flow. This transition highlights the importance of structural properties of the cupula in shaping the dynamic response of the vestibular system. The magnitude of the cupular deformation scales as $\hat{\eta}\sim a^{2}R\Omega_{0}/(\nu\kappa)=aR^{2}\Omega_{0}\rho/(E\mathcal{T})$ in this regime.
  • High-frequency regime, $\mathrm{St}\gg\max\{\kappa,1\}$: if the forcing frequency increases substantially inertial effects in the fluid must be considered and the deformation follows the angular displacement of the head. This type of motion could be expected when e.g. driving over an uneven surface at speed, or in a collision. In this regime the magnitude of the deformation is $\hat{\eta}\sim R^{2}\Omega_{0}\mu/(Ea)$.
In this work the cupula was modelled as an elastic solid of finite thickness, $t_{h}>0$, in contrast to previous analytical models that either treated it as a plate of vanishing thickness (Rabbitt & Damiano, 1992) or that did not include the cupula as a physical solid structure (Obrist, 2008). Experimental observations indicate that the cupula is not thin (Selva et al., 2009), motivating the inclusion of thickness as a parameter to more accurately capture its deformation in an analytically tractable framework. Our theoretical model suggests that:
  • For small thickness ratios $\beta=t_{h}/a$, the deformation resembles that of a clamped plate (a quartic profile). In this case, the cupular deformation scales as $\beta^{-3}$ for $\beta\ll 1$.
  • Increasing $\beta$ leads to a transition towards a quadratic deformation that matches the radially symmetric leading-order velocity profile $w_{0}$. In this case the deformation is smaller, scaling with $\beta^{-1}$.
@@
To verify the analytical findings, we conducted numerical simulations of the reduced equations using COMSOL Multiphysics. The numerical results showed excellent agreement with the asymptotic predictions, confirming the validity of the analytical approximations across a wide parameter space. Importantly, the numerical approach enabled us to also account for the influence of complex fluid–solid interaction boundary conditions and the nonlinear, advection, term in the Navier-Stokes equations that are otherwise intractable analytically, and comparison with the analytical solution points to their influence being small (and $\mathrm{Re}$ not being a relevant parameter in the problem).
@@
When the inertial terms were incorporated into the governing equations via inclusion of a finite Stokes number, $\mathrm{St}$, we observed important modifications to the system’s behaviour. For small Stokes numbers, and in studying the relaxation of the cupula to an initial deformation, the system exhibited overdamped dynamics. This behaviour is also predicted when inertia is neglected, and is consistent with the low-Reynolds-number assumption inherent to the vestibular fluid dynamics. However, for sufficiently large Stokes numbers, the system exhibited underdamped oscillations, with the cupular deformation following angular displacement even in the soft cupula regime for sufficiently high $\mathrm{St}$. This transition highlights the interplay between inertial and viscous forces in shaping the dynamic response of the system. Physiologically, this finding suggests that under certain conditions, such as during rapid head movements (e.g. during an impact), the vestibular system may exhibit enhanced sensitivity to displacement due to inertial effects.
@@
In each of the scenarios considered, our analytical approach enabled us to derive explicit expressions for the transition point between physical regimes, that is we obtained formulas for the critical values of the relevant system parameters at which the transition between different regimes occurs. These formulas lend insight into the fine balance between different components of the system, and enable us to speculate on how the vestibular system may have been fine tuned by evolution in different organisms, and/or key considerations in engineering an artificial vestibular system. We turn to such considerations next.